# Non-abelian Rees construction and pure motives

### v1

@@ 1.4. Functoriality
On the other hand, let $\mathcal{F}$ be the category of pairs $(G,{\rm X})$, where $G$ is a proreductive group over $K$ and $\rm X$} a fix-pointed affine $G$-scheme, and morphisms are equivariant maps.
@@
The functoriality of O’Sullivan’s construction ${\mathcal{C}}\mapsto{\rm{X}}_{\mathcal{C}}$ is then given by the mate101010cf. $\rm{nLab}$: mates, 1.1 (https://ncatlab.org/); in the scheme-theoretic context, this is just the base change morphism of SGA IV.
@@
in the $2$-category of essentially small tensor categories with lax tensor functors:
@@
$\Phi$ yields a map ${\mathcal{F}}({{\rm X}_{\mathcal{C}}^{\prime}},{\rm X})\to{\mathcal{E}}({\rm{Vec}}_{G}({\rm X}),{\mathcal{C}}^{\prime})$, functorial in $\rm X$} and $\mathcal{C}$$^{\prime}$, which admits a section by the previous lemma. It remains to show that $\Phi$ is faithful. In this generality, this requires more advanced tools, namely (part of) the recognition theorem for quotient stacks as in Savin [43] or Tonini [47]. Let us consider the quotient stacks $\mathfrak{X}=[{\rm X}/G],\mathfrak{X}^{\prime}=[{\rm X}^{\prime}/G^{\prime}]$.
These authors use a different notion of fiber functors: they are not necessarily faithful, but are right-exact in a technical sense (which uses the resolution property: any coherent module over $\mathfrak{X}$ is a quotient of an object of ${\rm{Vec}}\,\mathfrak{X}$ (${\rm{Vec}}_{G}({\rm X})$ viewed as a stack)111111this fact seems to have been first observed by Thomason [46].. Denoting by $Fib^{\prime}$ the 2-category of fiber functors in their sense (roughly speaking, tensor functors toward ${\rm{Vec}}_{K}$ which ‘preserve surjections’), they prove that the covariant composition
@@
The faithfulness of $\mathfrak{X}\mapsto Fib^{\prime}({\rm{Vec}}\,\mathfrak{X})$ together with the following lemma (valid without assuming that $\rm X$} is fix-pointed or affine) then implies the theorem.
@@
The functor $(G,{\rm X})\mapsto({\mathfrak{X}}\to B_{G})$ is faithful121212cf. [19, 3.2] in the case of homogeneous spaces, where an explicit quasi-inverse of the functor is produced; in this case, $\mathfrak{X}$} turns out to be a gerb - an incarnation of the Springer gerb..
Let $f,g$ be two equivariant maps ${\rm X^{\prime}}\to{\rm X}$ which induce the same commutative diagram (1.4). To an $S$-point $P^{\prime}$ of $\mathfrak{X}^{\prime}$, i.e. a $G^{\prime}_{S}$-torsor together with an equivariant map $P^{\prime}\to{\rm X}$, one gets equal $S$-points of $\mathfrak{X}$, therefore equal $S$-torsors $G\wedge_{G^{\prime}}f^{\ast}P^{\prime}=G\wedge_{G^{\prime}}g^{\ast}P^{\prime}$ with the same map to $X$ (via $f$ and $g$ respectively). Applying $1\wedge_{G}$-, one gets $f=g$. $\square$
@@
Morphisms of geometric stacks induce tensor functors between associated categories of vector bundles which have the property of “preserving surjections", a condition which is not intrinsic but depends on the ambient category of quasicoherent sheaves [11, §3.1, Ex. 3.5]. This property is not automatic: taking again Example 1.3.1, the composed functor $\phi:{\rm{Vec}}\,(\mathbb{P}^{1})\stackrel{{\scriptstyle\sim}}{{\to}}{\rm{Vec}}([\mathbb{A}^{2}/\mathbb{G}_{m}])\to{\rm{Vec}}([o/\mathbb{G}_{m}])={\rm{GrVec}}_{K}\to{\rm{Vec}}_{K},$ which is neither in $Fib^{\prime}\,{\mathcal{C}}$ nor in $Fib\,{\mathcal{C}}$, does not come by taking the fiber at any point of $\mathbb{P}^{1}$, but does come from taking the fiber at the origin $o$ of $\mathbb{A}^{2}$. In fact, the standard map ${\mathcal{O}}_{\mathbb{P}^{1}}^{2}\to{\mathcal{O}}_{\mathbb{P}^{1}}(1)$ defining $\mathbb{P}^{1}$}, which $\phi$ sends to 0, is surjective in the category of quasicoherent sheaves on $\mathbb{P}^{1}$, but not in the category of quasicoherent sheaves on the stack $\mathbb{A}^{2}/\mathbb{G}_{m}]$.
@@ 2. The restricted non-abelian Rees construction
The restricted non-abelian Rees construction deals with integral even $K$-tensor categories which satisfy further ‘finiteness’ conditions, and relate them to fix-pointed quasi-homogeneous affine varieties with respect to the action of a reductive group over $K$. The situation thus becomes closer to the original Rees construction, and amenable to computations.
@@ 2.1. A glimpse of the theory of tannakian hulls
Let $\mathcal{C}$} be a pseudo-abelian rigid $K$-tensor category with ${{\mathcal{C}}}({\bf 1},{\bf 1})=K$. If there is a fiber functor with coefficients in some extension $K^{\prime}$ of $K$, then $\mathcal{C}$} is integral even. O’Sullivan has shown that the converse is true.
@@
O’Sullivan [38]131313actually, O’Sullivan’s theory is more general: it deals with super fiber functor with values in any extension $K^{\prime}$ and yields super tannakian hulls, and not just tannakian hulls in the even case. and Coulembier [14] (by different means) have shown that $\omega$ factors through a neutral tannakian category ${\mathcal{C}}_{tann}$ over $K$ (the tannakian hull of $\mathcal{C}$}) which is universal in the sense that any fiber functor with coefficients in some $K$-extension factors through ${\mathcal{C}}_{tann}$.
The fiber functor $\omega$ induces a tensor equivalence ${\mathcal{C}}_{tann}\cong{\rm{Rep}}\,H,$ with $H={\rm{Aut}}^{\otimes}\omega_{\mid{\mathcal{C}}_{tann}}$. If $({\mathcal{C}},\bar{\omega},\sigma)\in{\mathcal{E}}$ is such that $\bar{\omega}=\omega\sigma$, then $H$ becomes a closed subgroup of $G={\rm{Aut}}^{\otimes}\bar{\omega}$.
The universality of ${\mathcal{C}}_{tann}$ translates into an equality of gerbs $Fib\,{\mathcal{C}}=Fib\,{\mathcal{C}}_{tann},$ which is equivalent to the classifying gerb $B_{H}$.
@@ 2.1.2. Examples
Any faithful tensor functor $\phi:{\mathcal{C}}\to{\mathcal{C}}^{\prime}$ induces a faithful exact tensor functor $\phi_{tann}:{\mathcal{C}}_{tann}\to{\mathcal{C}}^{\prime}_{tann}$, whence a commutative diagram
@@
Caveat. If $\phi$ is full, $\phi_{tann}$ need not be full; in other words, if $G$ is a quotient of $G^{\prime}$, $H$ need not be the image of $H^{\prime}$ in $G$. This is a crucial issue in this theory, as we shall see later 4.4.3, especially in the motivic context.
@@
Barbieri-Viale and Kahn [9] have defined a more general notion of ‘abelian tensor hull’ $T({\mathcal{C}})$ of $\mathcal{C}$}, with a similar universal property but with respect to any (possibly non-faithful) tensor functor with abelian tensor target (which can be taken to be $Vec_{K^{\prime}}$ if $\mathcal{C}$} is even - in which case $T({\mathcal{C}})$ is even). There are tensor functors ${\mathcal{C}}\to T({\mathcal{C}})$ (which is faithful if $\mathcal{C}$} is integral), and $T({\mathcal{C}})\to{\mathcal{C}}_{tann},\,T({\mathcal{C}})\to\bar{\mathcal{C}}$ (which are full).
@@ 2.2. Finiteness conditions
Condition $P3)$ holds for $\bar{\mathcal{C}}$ under $P2)$: by Weyl’s polarization-restitution theorem, ${\rm{Ind}}\,\bar{\mathcal{C}}({\bf 1},Sym({\mathcal{V}}^{m}\oplus({\mathcal{V}}^{\vee})^{n}))=(Sym(T^{m,n}V))^{G}$, for $m,n\geq d=\dim{\mathcal{V}}$, is generated by ${\rm{Ind}}\,\bar{\mathcal{C}}({\bf 1},Sym({\mathcal{V}}^{d}\oplus({\mathcal{V}}^{\vee})^{d}))=(Sym(T^{d,d}V))^{G}$ [27, §4], and the latter is a finitely generated algebra (Hilbert).
@@
If ${\mathcal{C}}={\rm{Vec}}_{G}({\rm X})$, for $G$ a reductive $K$-group and $X$ a quasi-homogeneous fix-pointed affine $G$-variety such that ${\rm{X}}_{0}(K)\neq\emptyset$, then conditions $P1,P2,P3$ hold.
@@ 2.3. Non-abelian Rees construction for integral even tensor categories
Note that a fix-pointed quasi-homogeneous variety $(G,{\rm X})$ with ${\rm X}_{0}(K)\neq\emptyset$ is the same thing as a $G$-variety $\rm X$ which admits a dominant equivariant map $G\to\rm X$. The image of this map is a homogeneous variety ${\rm X}_{0}=G/H$.
@@
2) The construction is functorial and provides an antiequivalence between the full subcategory of $\mathcal{E}$} of objects $\mathcal{C}$} satisfying $P1,P2,P3$, and the full subcategory of $\mathcal{F}$} of fix-pointed affine quasi-homogeneous varieties $(G,{\rm X})$ with ${\rm X}_{0}(K)\neq\emptyset$.
@@
$1)$ Under condition $P1$, we may and shall take $\bar{\omega}=\omega\sigma$.
To see that $\rm X$} is of finite type under P2 and P3, one notes that $\rm X$} is a limit of affine $G$-varieties ${\rm X}_{\lambda}$ of finite type [38, 9.6], and that the tower of categories $Vec_{G}({\rm X}_{\lambda})$ stabilizes under P2 and P3.
Then $\rm X$} is a quasi-homogeneous by P1, cf [38, 9.8, 9.13].
It remains to see that ${\rm X}_{0}(K)\neq\emptyset$. One has $Fib\,{\rm{Vec}}_{G}({\rm X_{0}})\cong B_{H}\cong\mathfrak{X}_{0}$. In the dictionary between gerbs over $B_{G}$ and homogeneous spaces [19, 3.2], $B_{H}\cong\mathfrak{X}_{0}$ corresponds to ${\rm X_{0}}\cong G/H$.
2) $(G,{\rm X})\mapsto({\mathcal{C}}=Vec_{G}({\rm X}),\overline{\omega})$ is fully faithful, as we have already seen. Let us give another proof of faithfulness, which does not rely on the recognition theorem for quotient stacks, but only on the familiar recognition for quotient gerbs (the usual tannakian correspondence).
Let $f,g$ induce the same functor $f^{\ast}=g^{\ast}$ on ${\rm{Vec}}\,_{G}{\rm X}$, hence on $Fib\,{\rm{Vec}}_{G}({\rm X})=Fib\,{\rm{Vec}}_{G}({\rm X_{0}})$. The idea is that ${\rm{Vec}}\,\mathfrak{X}_{0}$ being tannakian, $Fib\,{\rm{Vec}}\,\mathfrak{X}_{0}=Fib\,{\rm{Vec}}\,\mathfrak{X}$ is then equivalent to $\mathfrak{X}_{0}$, and one can copy Lemma 1.4.3, with a caveat: $\mathfrak{X}^{\prime}_{0}$ goes to $\mathfrak{X}_{0}$ when $f^{\ast}=g^{\ast}$ is faithful, but not in general. Note however that on replacing ${\mathcal{O}}({\rm X})$ by the image of ${\mathcal{O}}({\rm X}^{\prime})$ (which is also $G$-integral), one gets a $G$-variety $\rm Y$ which has the same properties as $\rm X$. One may thus assume $\rm Y=X$.
By usual tannakian theory (together with the fact that fiber functors in our sense are exact [38][24]), $Fib\,{\rm{Vec}}\,{\mathfrak{X}}_{0}\cong{\mathfrak{X}}_{0}$. And one concludes as before. (Note also that in this situation $Fib\,{\rm{Vec}}\,\mathfrak{X}\subset Fib^{\prime}\,{\rm{Vec}}\,\mathfrak{X}$ via the embedding $\mathfrak{X}_{0}\subset\mathfrak{X}$).

### v2

@@ 1.4. Functoriality
We do not assume any compatibility between $(\phi,\bar{\phi})$ and the projection functors $(\pi,\pi^{\prime})$; that is to say, we do not assume that $\phi$ preserves tensor radicals.
@@
Let $\mathcal{F}$ be the category of pairs $(G,{\rm X})$, where $G$ is a proreductive group over $K$ and $\rm X$} a fix-pointed affine $G$-scheme, and morphisms are equivariant maps.
@@
The functoriality of O’Sullivan’s construction ${\mathcal{C}}\mapsto{\rm{X}}_{\mathcal{C}}$ is then given by the mate
@@
in the $2$-category of essentially small tensor categories with lax tensor functors (cf. 1.2.4): namely:
@@
$\Phi$ yields a map ${\mathcal{F}}({{\rm X}_{\mathcal{C}}^{\prime}},{\rm X})\to{\mathcal{E}}({\rm{Vec}}_{G}({\rm X}),{\mathcal{C}}^{\prime})$, functorial in $\rm X$} and $\mathcal{C}$$^{\prime}$, which admits a section by the previous lemma. Therefore $\Phi$ is full and essentially surjective, and it remains to show that $\Phi$ is faithful.
In this generality, faithfulness requires more advanced tools, namely (part of) the recognition theorem for quotient stacks as in Savin [45] or Tonini [49]. Let us consider the quotient stacks
$\mathfrak{X}=[{\rm X}/G],\;\mathfrak{X}^{\prime}=[{\rm X}^{\prime}/G^{\prime}].$
and identify ${\rm{Vec}}\,\mathfrak{X}$ with ${\rm{Vec}}_{G}({\rm X})$ as usual.
These authors use a different notion of fiber functors as here (cf. 1.2.1): their fiber functors are not necessarily faithful, but are right-exact in a technical sense: roughly speaking, they are tensor functors toward ${\rm{Vec}}_{K}$ that ‘preserve surjections’ (this implicitly uses the resolution property: any coherent module over $\mathfrak{X}$ is a quotient of an object of ${\rm{Vec}}\,\mathfrak{X}$, a fact first observed by Thomason [48]). Let us denote by $Fib^{\prime}$ the 2-category of fiber functors in their sense. They prove that the composition of contravariant functors
@@
The faithfulness of $\mathfrak{X}\mapsto Fib^{\prime}({\rm{Vec}}\,\mathfrak{X})$ together with the following lemma (also valid without assuming that $\rm X$} is fix-pointed or affine) then implies the theorem.
@@
The functor $(G,{\rm X})\mapsto{\mathfrak{X}}$ is faithful101010cf. [20, 3.2] in the case of homogeneous spaces - where the functor $(G,{\rm X})\mapsto({\mathfrak{X}}\to B_{G})$ is an equivalence; an explicit quasi-inverse is made explicit. In this case, $\mathfrak{X}$} turns out to be a gerb - an incarnation of the Springer gerb..
Let $f,g$ be two equivariant maps ${\rm X^{\prime}}\to{\rm X}$ that induce the same map ${\mathfrak{X}}^{\prime}\to{\mathfrak{X}}$. Then they induce the same commutative diagram (1.4). To an $S$-point $P^{\prime}$ of $\mathfrak{X}^{\prime}$, i.e. a $G^{\prime}_{S}$-torsor together with an equivariant map $P^{\prime}\to{\rm X}$, one gets equal $S$-points of $\mathfrak{X}$, therefore equal $G_{S}$-torsors $G\wedge_{G^{\prime}}f^{\ast}P^{\prime}=G\wedge_{G^{\prime}}g^{\ast}P^{\prime}$ with the same map to $X$ (via $f$ and $g$ respectively). Applying $1\wedge_{G}$- and letting $P^{\prime}$ vary, one gets $f=g$. $\square$
@@
Morphisms of geometric stacks induce tensor functors between associated categories of vector bundles that have the property of ‘preserving surjections’, a condition that is not intrinsic but depends on the ambient category of quasicoherent sheaves [11, §3.1, Ex. 3.5]. This property is not automatic: taking again Example 1.3.1, the composed functor $\phi:{\rm{Vec}}\,(\mathbb{P}^{1})\stackrel{{\scriptstyle\sim}}{{\to}}{\rm{Vec}}([\mathbb{A}^{2}/\mathbb{G}_{m}])\to{\rm{Vec}}([o/\mathbb{G}_{m}])={\rm{GrVec}}_{K}\to{\rm{Vec}}_{K},$ which is neither in $Fib^{\prime}\,{\mathcal{C}}$ nor in $Fib\,{\mathcal{C}}$, does not come from taking the fiber at any $K$-point of $\mathbb{P}^{1}$ but does come from taking the fiber at the origin $o$ of $\mathbb{A}^{2}$: indeed, the standard map ${\mathcal{O}}_{\mathbb{P}^{1}}^{2}\to{\mathcal{O}}_{\mathbb{P}^{1}}(1)$ defining $\mathbb{P}^{1}$}, which $\phi$ sends to 0, is surjective in the category of quasicoherent sheaves on $\mathbb{P}^{1}$, but not in the category of quasicoherent sheaves on the stack $\mathbb{A}^{2}/\mathbb{G}_{m}]$.
@@ 2. The restricted non-abelian Rees construction
The restricted non-abelian Rees construction deals with integral even $K$-tensor categories that satisfy further ‘finiteness’ conditions, and relate them to fix-pointed affine quasi-homogeneous varieties with respect to the action of a reductive group over $K$. The situation thus becomes closer to the original Rees construction, and amenable to computations.
@@ 2.1. A glimpse of the theory of tannakian hulls
Let $\mathcal{C}$} be a pseudo-abelian rigid $K$-tensor category with ${{\mathcal{C}}}({\bf 1},{\bf 1})=K$. If there is a fiber functor with coefficients in some extension $K^{\prime}$ of $K$, then $\mathcal{C}$} is integral even. O’Sullivan has shown that the converse is true [40, Th. 10.10] (specialized to the even case).
@@
O’Sullivan [40]111111actually, O’Sullivan’s theory is more general: it deals with super fiber functor with values in any extension $K^{\prime}$ and yields super tannakian hulls, and not just tannakian hulls in the even case. and Coulembier [14] (by different means) have shown that $\omega$ factors through a neutral tannakian category ${\mathcal{C}}_{tann}$ over $K$ (the tannakian hull of $\mathcal{C}$}), which is universal in the sense that any fiber functor with coefficients in some $K$-extension factors through a fiber functor
$\omega_{tann}:{\mathcal{C}}_{tann}\to{\rm{Vec}}_{K}.$
Moreover, any object of ${\mathcal{C}}_{tann}$ is a quotient (or, dually, a subobject) of an object of $\mathcal{C}$} [40, p.2].
The fiber functor $\omega$ induces a tensor equivalence ${\mathcal{C}}_{tann}\cong{\rm{Rep}}\,H,$ with $H={\rm{Aut}}^{\otimes}\omega_{tann}$. If $({\mathcal{C}},\bar{\omega},\sigma)\in{\mathcal{E}}$ is such that $\bar{\omega}=\omega\sigma$, then $H$ becomes a closed subgroup of $G={\rm{Aut}}^{\otimes}\bar{\omega}$ by the criterion [19, 2.21].
The universality of ${\mathcal{C}}_{tann}$ translates into an equality of gerbs $Fib\,{\mathcal{C}}=Fib\;{\mathcal{C}}_{tann},$ the latter being equivalent to the classifying gerb $B_{H}$.
@@ 2.1.2. Examples
Any faithful tensor functor $\phi:{\mathcal{C}}\to{\mathcal{C}}^{\prime}$ induces a faithful exact tensor functor $\phi_{tann}:{\mathcal{C}}_{tann}\to{\mathcal{C}}^{\prime}_{tann}$, whence (via the fiber functor $\omega^{\prime}_{tann}$ and tannakian duality) a commutative diagram
@@
Caveat. If $\phi$ is full, $\phi_{tann}$ need not be full; in other words, if $G$ is a quotient of $G^{\prime}$, $H$ need not be the image of $H^{\prime}$ in $G$. This is a crucial issue in this theory, as we shall see later (4.4.2), especially in the motivic context.
@@
Barbieri-Viale and Kahn [9] have defined a more general notion of ‘abelian tensor hull’ $T({\mathcal{C}})$ of $\mathcal{C}$}, with a similar universal property but with respect to any, possibly non-faithful, tensor functor with abelian tensor target $\mathcal{C}$$^{\prime}$ (which can be taken to be $Vec_{K^{\prime}}$ if $\mathcal{C}$} is even - in which case $T({\mathcal{C}})$ is even). There are tensor functors ${\mathcal{C}}\to T({\mathcal{C}})$ (faithful if $\mathcal{C}$} is integral), and $T({\mathcal{C}})\to{\mathcal{C}}_{tann},\,T({\mathcal{C}})\to\bar{\mathcal{C}}$ (both full).
@@ 2.2. Finiteness conditions
Condition $P3)$ holds for $\bar{\mathcal{C}}$ under $P2)$: by Weyl’s polarization-restitution theorem, ${\rm{Ind}}\,\bar{\mathcal{C}}({\bf 1},Sym({\mathcal{V}}^{m}\oplus({\mathcal{V}}^{\vee})^{n}))=(Sym(T^{m,n}V))^{G}$, for $m,n\geq d=\dim{\mathcal{V}}$, is generated by ${\rm{Ind}}\,\bar{\mathcal{C}}({\bf 1},Sym({\mathcal{V}}^{d}\oplus({\mathcal{V}}^{\vee})^{d}))=(Sym(T^{d,d}V))^{G}$ [28, §4], and the latter is a finitely generated algebra (Hilbert).
@@
If ${\mathcal{C}}={\rm{Vec}}_{G}({\rm X})$, for $G$ a reductive $K$-group and $X$ a quasi-homogeneous fix-pointed affine $G$-variety such that the open orbit ${\rm{X}}_{0}$ has a $K$-rational point, then conditions $P1,P2,P3$ hold.
@@ 2.3. Non-abelian Rees construction for integral even tensor categories
Note that a quasi-homogeneous variety $(G,{\rm X})$ with a $K$-point in the open orbit ${\rm X}_{0}=G/H$ is the same thing as a $G$-variety $\rm X$ that admits a dominant equivariant map $G\to\rm X$.
@@
2) The construction is functorial. With the notation of §1.4.2, it provides an antiequivalence between the full subcategory of $\mathcal{E}$} of objects $\mathcal{C}$} satisfying $P1,P2,P3$, and the full subcategory of $\mathcal{F}$} of fix-pointed affine quasi-homogeneous varieties $(G,{\rm X})$ with ${\rm X}_{0}(K)\neq\emptyset$.
@@
$1)$ Under condition P1, $\mathcal{C}$} is even, and $\bar{\mathcal{C}}$ is neutral: we may and shall take $\bar{\omega}=\omega\sigma$.
To see that $\rm X$} is of finite type under P2 and P3, one notes that $\rm X$} is a limit of affine $G$-varieties ${\rm X}_{\lambda}$ of finite type [40, 9.6], and that the tower of categories $Vec_{G}({\rm X}_{\lambda})$ stabilizes under P2 and P3.
Then $\rm X$} is a quasi-homogeneous by P1, cf [40, 9.7, 9.13]; more precisely, ${\rm X}_{0}$ is schematically dense in $\rm X$ (hence $\rm X$ is reduced).
It remains to see that ${\rm X}_{0}(K)\neq\emptyset$. One has $Fib\,{\rm{Vec}}_{G}({\rm X_{0}})\cong B_{H}\cong\mathfrak{X}_{0}$. In the dictionary between gerbs over $B_{G}$ and homogeneous spaces [20, 3.2], $B_{H}\cong\mathfrak{X}_{0}$ corresponds to ${\rm X_{0}}\cong G/H$.
2) $(G,{\rm X})\mapsto({\mathcal{C}}=Vec_{G}({\rm X}),\overline{\omega})$ is fully faithful, as we have already seen. Let us give another proof of faithfulness, which does not rely on the recognition theorem for quotient stacks, but only on the familiar recognition theorem for quotient gerbs (i.e. the usual tannakian correspondence).
Let $f,g$ induce the same functor $f^{\ast}=g^{\ast}$ on ${\rm{Vec}}_{G}({\rm X})$, hence on $Fib({\rm{Vec}}\,{\mathfrak{X}})=Fib({\rm{Vec}}\,{\mathfrak{X}_{0}})$. Since ${\rm{Vec}}\,\mathfrak{X}_{0}$ is tannakian, $Fib({\rm{Vec}}\,\mathfrak{X}_{0})$ is equivalent to $\mathfrak{X}_{0}$, and one can copy Lemma 1.4.3, with a caveat: $\mathfrak{X}^{\prime}_{0}$ goes to $\mathfrak{X}_{0}$ when $f^{\ast}=g^{\ast}$ is faithful, but not in general. Note however that on replacing ${\mathcal{O}}({\rm X})$ by the image of ${\mathcal{O}}({\rm X}^{\prime})$ (which is also $G$-integral) under $f^{\ast}=g^{\ast}$, one gets a $G$-variety $\rm Y$ that has the same properties as $\rm X$. One may thus assume $\rm Y=X$.
By usual tannakian theory (together with the fact that faithful tensor functors on a tannakian category are exact [40][25]), $Fib({\rm{Vec}}\,{\mathfrak{X}}_{0})\cong{\mathfrak{X}}_{0}$, and one concludes as in the proof of Theorem 1.4.2. (Note also that in this situation $Fib({\rm{Vec}}\,\mathfrak{X})\subset Fib^{\prime}({\rm{Vec}}\,\mathfrak{X})$ via the embedding $\mathfrak{X}_{0}\subset\mathfrak{X}$).